you need 500mL of 1.5 M HCL for an experiment but the stock room only has 12 M HCl how do you make 1.5 M HCL




V1= (c2)(v2)/c1mole/ liter = (1.5m)(500mL)/12 mole/liter = 41.6?


This is the answer I got but it doesn't sound right consider its significantly higher than 1.5
Answer #1

Your process is correct.

1st we need to calculate the volume of stock solution required.

V1 = C2*V2/C1 = 1.5M*500 ml / 12 M= 62.5 mL

So we need 62.4 mL of stock solution.

Take 62.4 mL of this stock solution.Add enough water to it tomake the volume to 500 mL